OpenStudy (jgirl128):

Weird question, I guess. This is my question: 1. Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

OpenStudy (jgirl128):

and this is what I got: $500 = [0.5(1.4) – 0.5)]f It costs them $0.50 to purchase each flower. Then they sell the flowers at the purchase price plus 40% so $0.50 + 40%. So the cost they spent on all the flowers is 0.50x (Use c to represent this) So this would mean they sell each flower for $0.70. (Use s to represent this) So the profit (use p to represent this) is the difference between the cost and sold. p = s - c p = 0.70x - 0.50x p = (0.70 - 0.50)x p = (0.20)x p = 0.20x So the profit made on each flower sold is $0.20. To make a profit of $500... 500 = 0.20x Once I solved for x I got x = 2500 So you must buy and sell 2,500 flowers

OpenStudy (jgirl128):

and my teacher responded back to me and told me to use this equation: $500 = [0.5(1.4) – 0.5)]f so I'm confused

OpenStudy (tkhunny):

Generally, it helps if you provide some clue as to where it is you might be stuck. Cost: 0.50 Retail Price: 0.50 * 1.40 = 0.70 Profit: 0.70 - 0.50 = 0.20 You can see all the pieces, in there.

OpenStudy (jgirl128):

I'll try to figure it out. thanks for the help! :)